OpenStudy (anonymous):

Jacob spends 60 minutes in the gym every day doing freehand exercises and running on the treadmill. He spends 30 minutes more running on the treadmill than doing freehand exercises. Part A: Write a pair of linear equations to show the relationship between the number of minutes Jacob does freehand exercises (y) and the number of minutes he runs on the treadmill(x). (5 points) Part B: How much time does Jacob spend on doing freehand exercises? (3 points) Part C: Is it possible for Jacob to have spent 40 minutes running on the treadmill? Explain your reasoning. (2 points)

OpenStudy (anonymous):

I need help w/ Part C.

OpenStudy (anonymous):

Uhm well. Yes it is possible for Jacob to have spent 40 minutes on the treadmill. But I'm not understanding what part c is asking?

OpenStudy (anonymous):

Hold one, here's what I got for the answers: Part A.) A pair of linear equations would be: x+ y = 60, this would be because he spent 60 minutes doing both of his exercises. x = y + 30 Part B.) He spent 15 minutes doing freehand exercises.

OpenStudy (anonymous):

Ok. This looks like in part C you would be using the equation x + y = 60. So if Jacob spent 15 minutes doing free hand exercises you would put that in the equation as x. Solve for y. So you would subtract 15 minutes from both sides of the equation getting y = 60 - 15 so y = 45. So for part be I would say. Yes it is possible for Jacob to spend 40 minutes on the treadmill, and then you would write out why. Do you need further help? I would be happy to help!

OpenStudy (anonymous):

Thank you so much, you're very generous! Yeah, I could use some help w/ another question, would you mind helping me? :)

OpenStudy (anonymous):

Your welcome! I would be obliged to help you further! What other questions do you have?

OpenStudy (anonymous):

Alright before I send you another question, I'm still a little confused as to why he can spend 40 minutes on the treadmill.

OpenStudy (anonymous):

@JoMary where'd you go? :-)

OpenStudy (anonymous):

Ok. It is possible for Jacob to have spent 40 minutes running on the treadmill, because say Jabob took a 5 minute break in between his exercises. So if he spent 60 minutes doing both his exercises and 15 minutes doing free hand exercises, he would've only got 35 minutes of exercises on free hand and any other exercises he did, therefore if he exercises for 40 minutes after his other exercises he would be at a total of 60 minutes of exercising. Does this help? I just went to check on another question...thats all. But I'm here now! =)

OpenStudy (anonymous):

I'm just rereading your answer, and seeing why the answer is the answer. :)

OpenStudy (anonymous):

Yes, it really helps, alright, ready for the 2nd question? :)

OpenStudy (anonymous):

Yes I am ready! xD

OpenStudy (anonymous):

Alright: A system of equations is shown below. 3x + 8y = 12 2x + 2y = 3 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

Do you have any parts figured out?

OpenStudy (anonymous):

No...I don't really understand it that much. Somewhat, not really.

OpenStudy (anonymous):

Ok. Give me a minute.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Looks like I need to use the ellimination method. -2/3 x 3x(-2)+8y(-2) = 121 over 2x(3) + 2y(3) = 3(3) = 9 3x + 8y = 12 2x + 2y = 3 -6x -16y = -24 6x + 6y = 9 -10y = -15 /-10 /-10 y = 1.5 Sub in y into the equation and solve for x I was alittle confused for a moment, but thats your answer. Does this help?

OpenStudy (anonymous):

Hold on one second. :)

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

I'm confused as to how you got: -2/3 x 3x(-2)+8y(-2) = 121 over

OpenStudy (anonymous):

That is the ellimination method. Haven't you learned that in previous lessons?

OpenStudy (anonymous):

Oh wait...never mind, sorry. I confused myself! And for Part B:, I'd just show how I checked my work?

OpenStudy (anonymous):

Alrighty. Its okay, I don't mind. Yes. For part b you would just show how you checked your work to make sure it was right. Do you have any further questions?

OpenStudy (anonymous):

Yes, a couple more, would you mind helping me? :)

OpenStudy (anonymous):

Actually, I meant to say 1 more, sorry. I was looking at the wrong thing.

OpenStudy (anonymous):

I would love to! :)

OpenStudy (anonymous):

Great! Part A: Explain why the x-coordinates of the points where the graphs of the equations \[y = 8^{x} and y = 2^{x+2} intersect are the solutions of the equation 8^{x} = 2^{x+2}. Part B: Make tables to find the solution to 8^{x} = 2^{x+2}. Take the integer values of x between -3 and 3. Part C: How can you solve the equation 8^{x} = 2^{x+2} graphically?

OpenStudy (anonymous):

I am sorry. But I have not covered this material yet. Do you have any other questions?

OpenStudy (anonymous):

Wait, is it you don't understand the inequalities, or you don't know how to solve it? :)

OpenStudy (anonymous):

I don't know how to solve it. What grade is this material? I know how to solve inequalites but not intense big ones like these....sorry.

OpenStudy (anonymous):

It's cool, it's 9th grade algebra.

OpenStudy (anonymous):

Thank you for the help :)

OpenStudy (anonymous):

Well I'm on in 7th grade. So...yea...this material is not known to me..yet. Your welcome. I am glad I could be of help for most all your questions.